OpenStudy (anonymous):

Larry has taken out a loan for college. He started paying off the loan with a first payment of $150. Each month he pays, he wants to pay back 1.3 times as the amount he paid the month before. Explain to Larry how to represent his first 15 payments in sigma notation. Then explain how to find the sum of his first 15 payments, using complete sentences. Explain why this series is convergent or divergent.

OpenStudy (anonymous):

@goformit100 @Kainui @adrynicoleb @anonymous_user @beccaboo333 @ByteMe @sourwing @sammixboo @tHe_FiZiCx99 @Cpt.Sparz @charliealbert @Drsuz98 @e.mccormick @faisalalif1999 @queenbee8 @iambatman @Luigi0210 @myininaya @nincompoop @nikosis01

OpenStudy (anonymous):

first==150 second- 150*1.3 third---150*1.3*1.3 it is geometric progression \[\sum_{1}^{15}a.r^(n-1)\]

OpenStudy (anonymous):

\[\sum_{1}^{15}a*r^{n-1}\]

OpenStudy (anonymous):

http://www.ilovemaths.com/3gpsum.asp

OpenStudy (anonymous):

So what do I write in sentences?

OpenStudy (anonymous):

First payment larry makes =150 second “ “ “ “=150*1.3 third............=150*1.3*1.3 similarly 15th payment=150*1.3^14 this series is in the form of GP

OpenStudy (anonymous):

@minoz Is the coverigent or the other way?

OpenStudy (anonymous):

a series is convergent if limit exists here as limit is 15 it is convergent

OpenStudy (anonymous):

and we get result as some number if result is infinity then series is divergent

OpenStudy (anonymous):

@minoz Daniel needs to study a possible link between anti-inflammatory drugs and kidney disease. Help Daniel design a study that would be appropriate. Explain the type of study, the sampling method, the parameter of interest, and the statistic he could use to estimate that parameter. What conclusions would Daniel be able to draw from this study?

OpenStudy (anonymous):

i dont have idea on sampling sorry

OpenStudy (anonymous):

Answer the attached

OpenStudy (anonymous):

@minoz

OpenStudy (e.mccormick):

Use the circumference of the small one to find the rotational distance. Then use the rotational distance to find the angular measure of the large. Same sort of thing for part two, but large to small.